# Provably Safe Generative Sampling with Constricting Barrier Functions

### v1

@@ Abstract
Flow-based generative models, such as diffusion models and flow matching models, have achieved remarkable success in learning complex data distributions. However, a critical gap remains for their deployment in safety-critical domains: the lack of formal guarantees that generated samples will satisfy hard constraints. We address this by proposing a safety filtering framework that acts as an online shield for any pre-trained generative model. Our key insight is to cooperate with the generative process rather than override it. We define a constricting safety tube that is relaxed at the initial noise distribution and progressively tightens to the target safe set at the final data distribution, mirroring the coarse-to-fine structure of the generative process itself. By characterizing this tube via Control Barrier Functions (CBFs), we synthesize a feedback control input through a convex Quadratic Program (QP) at each sampling step. As the tube is loosest when noise is high and intervention is cheapest in terms of control energy, most constraint enforcement occurs when it least disrupts the model’s learned structure. We prove that this mechanism guarantees safe sampling while minimizing the distributional shift from the original model at each sampling step, as quantified by the KL divergence. Our framework applies to any pre-trained flow-based generative scheme requiring no retraining or architectural modifications. We validate the approach across constrained image generation, physically-consistent trajectory sampling, and safe robotic manipulation policies, achieving 100% constraint satisfaction while preserving semantic fidelity.
@@ 1 Introduction
Refer to caption Unguided sampling process: $\mathrm{d}x=f_{\theta}(x,t)\mathrm{d}t+g(t)\mathrm{d}w$ Safe sampling process: $\mathrm{d}x=[f_{\theta}(x,t)+u]\mathrm{d}t+g(t)\mathrm{d}w$,  s.t. $x(t)\in\tilde{\mathcal{C}}(t)$ for all $t$
@@ 4 Guidance of flow-based models via constricting CBFs
Having established the sampling dynamics (1) and the CBF framework for safe control, we now address the central question: how can we enforce hard constraints on the output of a pre-trained generative model without modifying the model itself? The key observation is that by introducing a control input to the sampling SDE (1), we arrive at a control-affine structure for the control systems in Section 3.2. This additive control input $u$ can be designed to steer sampling trajectories toward the safe set while preserving the learned distribution. We frame this as a control synthesis problem where the guided sampling process is governed by:
@@ 4.1 CBFs for flow-based generative sampling
A fundamental challenge in generative sampling is that the initial prior noise $x(T)$ is drawn from a distribution that is easy to sample from, typically the standard Gaussian. This means that $x(T)$ almost certainly resides outside the target safe set $\mathcal{C}$. Traditional CBF formulations for static sets would require an immediate, high-energy intervention to push $x(t)$ close to $\mathcal{C}$, which would violate the model’s generative intent and produce samples that no longer represent the learned distribution. Our framework is motivated by the need for minimal interventions and a smooth approach towards $\mathcal{C}$. We achieve this by defining a constricting barrier $\tilde{h}(x,t)=h(x)+\epsilon(x(T),t)$. This is a time-varying barrier that constricts with the sampling process. We characterize safety through a constricting superlevel tube $\tilde{\mathcal{C}}(t)$ with respect to the constricting barrier $\tilde{h}$ as,
@@
We design $\tilde{\mathcal{C}}(t)$ such that initially, the set $\tilde{\mathcal{C}}(T)$ is sufficiently relaxed to contain the noise sample $x(T)$, and at the end of sampling, it recovers the target set $\mathcal{C}$, i.e., $\tilde{\mathcal{C}}(0)=\mathcal{C}$. Since the sampling process evolves in reverse time from $t=T$ to $t=0$, the safety tube $\tilde{\mathcal{C}}(t)$ must remain occupied by $x(t)$ as $t$ decreases, a property we term reverse invariance. As derived in Section 3.2, the reverse-time CBF condition requires bounding $\nabla\tilde{h}\cdot f$ from above. The following design of the constricting barrier function handles this through a time-varying control barrier function.
@@
Given a CBF $h:\mathbb{R}^{n}\to\mathbb{R}$ and a terminal condition $x(T)\in\mathbb{R}^{n}$, a constricting barrier function is $\tilde{h}(x,t)=h(x)+\epsilon(x(T),t)$, where $\epsilon$ is a $C^{1}$ function satisfying:
@@
Unlike other works that consider the problem of designing feedback controllers that enforce safety of stochastic dynamical systems Clark (2021), our approach assumes that the feedback controller has knowledge of the state as well as the noise. While this assumption is more restrictive, it allows us to obtain stronger guarantees. In particular, with our design approach, we can enforce invariance of the constricting safety tube exactly, whereas by So et al. (2023), invariance does not exactly or almost surely for the CBFs considered in Clark (2021).
@@
where the optimization is performed for the current state-noise pair $(x,\xi)$ at time $t$, yielding the control $u^{*}(x,\xi,t)$.
@@
$D_{\text{KL}}(p_{\text{safe}}(0)\parallel p(0))=\frac{1}{2}\mathbb{E}_{p_{\text{safe}}(0)}\left[\int_{0}^{T}\frac{\|u(x,t)\|^{2}}{g(t)^{2}}\mathrm{d}t\right].$ (9)
@@ 4.2 Discrete-time implementation
We note that the discrete-time constraint (11) is derived via a first-order Taylor series expansion of $\tilde{h}$ around $(x_{k},t_{k})$, and is therefore only a sufficient condition for $\tilde{h}(x_{k-1},t_{k-1})\geq 0$ only up to linearization error. For barrier functions with high curvature or large step sizes $\Delta t$, the true nonlinear barrier value could deviate from the linear prediction. In practice, the margin provided by the class-$\mathcal{K}$ term $\gamma(\tilde{h}_{k})\Delta t$ and the conservative initialization $\epsilon_{0}$ absorb this error. Across all experiments in Section 5, we observed no constraint violations at the final sample, empirically validating the linearization for the step sizes used. The discrete-time constraint (11) provides an approximate safety guarantee: the first-order Taylor expansion introduces an $O(\Delta t^{2})$ residual that is not formally bounded in our analysis. The additive margin $c>0$ can be introduced in the initialization $\epsilon_{0}$ and the class-$\mathcal{K}$ damping $\gamma$ to provide empirical robustness to this discretization error. A rigorous discrete-time CBF analysis that establishes explicit step-size-dependent bounds on constraint violation is an important direction for future work. Nonetheless, across all experiments with step sizes ranging from $\Delta t$ = 0.01 (Section 5.1) to $\Delta t=0.005$ (Section 5.3), we observed zero constraint violations at the final sample.
@@
1
2Initialize: Sample $x_{K}\sim p(T)$ (e.g., $x_{K}\sim\mathcal{N}(0,I)$), set $t_{K}=T$, $\Delta t=T/K$
3for $k=K,K-1,\ldots,1$ do
4   Sample noise: $\xi_{k}\sim\mathcal{N}(0,I)$
5  Compute noise-induced drift: $d_{\text{noise}}=g(t_{k})\sqrt{\Delta t}\,\xi_{k}$
6  Compute constricting barrier: $\tilde{h}_{k}=h(x_{k})+\epsilon(x_{K},t_{k})$
7  Solve safety-constrained QP:
$\min_{u_{k}}\quad$ $\frac{1}{2}\|u_{k}\|^{2}$
s.t. $\tilde{h}_{k}+\nabla_{x}\tilde{h}_{k}\cdot\left(-[f_{\theta}(x_{k},t_{k})+u_{k}]\Delta t+d_{\text{noise}}\right)-\frac{\partial\epsilon}{\partial t}(x_{K},t_{k})\Delta t\geq(1-\gamma(\tilde{h}_{k})\Delta t)\,\tilde{h}_{k}$
8  Apply control and update state: $x_{k-1}=x_{k}-f_{\theta}(x_{k},t_{k})\Delta t-u_{k}^{*}\Delta t+d_{\text{noise}}$
9  Update time: $t_{k-1}=t_{k}-\Delta t$
10
return $x_{0}$
@@ 5 Experiments
All experiments use the linear constriction $\epsilon(x(T),t)=\epsilon_{0}\cdot\frac{t}{T}$ where $\epsilon_{0}=\max(0,-h(x(T)))+c$, where $c>0$ is a small positive margin. This margin serves two purposes: it provides a buffer against the linearization error introduced by the discrete-time Taylor series expansion in (11), and it prevents the barrier from operating exactly on its zero level set at initialization, where numerical precision could cause spurious violations. In all experiments, we set $c=0.1$. We choose the class-$\mathcal{K}$ function $\gamma(h)=\alpha h$ with $\alpha=0.5$. The QPs in Algorithm 1 are solved using CVXPY with the OSQP solver. Our implementation is available at [anonymized for review]. Our experiments were conducted on an Intel i9-9900 machine with 128GB RAM and an Nvidia Quadro RTX 4000 GPU.
@@ 5.1 Physics-consistent trajectory generation for the Lorenz system
Figure 2(b) reveals the evolution of the safety tube. At $t=1$, the noise sample $x(T)\sim\mathcal{N}(0,I)$ has large cumulative physics error, requiring relaxation $\epsilon_{0}\approx 16$—over four orders of magnitude larger than our target tolerance $e=0.001$. The blue line shows the linear constriction $\epsilon(x(T),t)=\epsilon_{0}\cdot(t/T)$ to zero at $t=0$. The constriction barrier $\tilde{h}(x,t)$ (red) remains non-negative throughout, verifying reverse invariance (Theorem 4.1). Fluctuations during $t\approx 0.8$-$1.0$ reflect the high-noise regime where $g(t)$ is large. At $t=0$, $\tilde{h}(x,0)\approx 0$ with $\epsilon(x(T),0)=0$ implies $h(x(0))\approx 0$, confirming that our minimum-norm control maximally exploits the constraint tolerance.
Figure 2(c) shows control effort $\|u\|^{2}$ concentrated at sampling onset ($t\approx 0.8$-$1.0$), peaking around 130-140, then rapidly decaying to near-zero by $t\approx 0.6$. This front-loaded intervention exploits that control is cheaper when $g(t)$ is large. The minimal effort for $t<0.6$ indicates that the pre-trained model has implicitly learned the Lorenz system structure from training data, requiring only minor corrections to enforce hard guarantees.
@@ 5.2.2 Regional color intensity constraints
Figure 4: Regional color intensity constraints. All images enforce different RGB values in the lower third. (a) Moderate intensity: The target pixel color is black, and the spatial mask $v(\mathbf{p})$ goes from $0.5\to 0$, allowing the diffusion model to produce clean furniture. (b) Projection: Projection-based constraint enforcement (Zampini et al., 2025), constraint are same as in (a) with target pixel color black in the bottom one-third of the image. Projection at every sampling step causes intense constraint enforcement, which leads to loss of semantic information. (c) Low intensity: The target color is brown, and the spatial mask $v(\mathbf{p})$ goes from $0.2$ at the bottom to 0 at the lower-third boundary, producing a semantically meaningful image of a room with a brown carpet.
@@ 5.3 Smooth robot policy generation
where $e_{\text{smooth}}=1.5$ is the tolerance that bounds the maximum allowed jerk magnitude, and $\Delta_{s}$ is the time discretization of the environment. The safe set $\mathcal{C}=\{x\mid h(x)\geq 0\}$ encodes the set of all action sequences whose average control variation magnitude remains below the prescribed threshold $e_{\text{smooth}}$. We apply our CBF shield during the sampling process using Algorithm 1 with the same linear constriction scheme as in the previous experiments.
@@
Table 1: Quantitative comparison on Push-T with smoothness constraints over 100 episodes. Diffusion Policy with DDPM (DP) and DDIM (DP-DDIM) violate the smoothness constraint an average of 12 and 16 times per episode, respectively, whereas our CBF-guided framework achieves zero violations while maintaining identical task reward. The computational overhead of solving the QP (8) at each sampling step is modest relative to the environment control rate.
@@ 6 Limitations and future work
Several extensions follow naturally from our framework. First, extending the approach to latent diffusion models would enable scalability to higher-resolution generation. We investigated extending our framework to latent diffusion models (Rombach et al., 2022) where the diffusion process operates in a compressed latent space $z=E(x)$ and the constraint is defined in image space. While the QP structure is preserved, the latent gradient is computed via a single backward pass through the decoder. Our preliminary experiments with Stable Diffusion v1.5 revealed that exact constraint satisfaction does not transfer from latent to image space (see Appendix B.2). The VAE decoder is not a diffeomorphism: it is locally non-invertible and introduces reconstruction artifacts that distort the constraint boundary. As a result, the CBF shield successfully biases the latent trajectory toward the constrained region but cannot guarantee pixel-level fidelity in the decoded image. Achieving exact constraints in latent diffusion models likely requires either (i) a constraint-aware fine-tuning of the decoder to improve local invertibility near the constraint boundary, or (ii) a hybrid approach that applies coarse guidance in latent space and a final correction step in pixel space after decoding. Second, enforcing state-space safety constraints in robotic policies, such as collision avoidance, requires a dynamics model to map actions to states. Combining our framework with learned dynamics models or neural ODEs is a promising direction. Third, our current implementation solves one QP per sampling step using a general-purpose solver. Exploiting the closed-form solution of the single-constraint QP or using an MPC framework across many steps could significantly reduce computational overhead and obtain bounds on the KL divergence over the entire sampling horizon.

### v2

# arXiv:2602.21429v2[cs.LG] 27 Feb 2026

@@ Abstract
Flow-based generative models, such as diffusion models and flow matching models, have achieved remarkable success in learning complex data distributions. However, a critical gap remains for their deployment in safety-critical domains: the lack of formal guarantees that generated samples will satisfy hard constraints. We propose a safety filtering framework that acts as an online shield for any pre-trained generative model. Our key insight is to cooperate with the generative process rather than override it. We define a constricting safety tube that is relaxed at the initial noise distribution and progressively tightens to the target safe set at the final data distribution, mirroring the coarse-to-fine structure of the generative process itself. By characterizing this tube via Control Barrier Functions (CBFs), we synthesize a feedback control input through a convex Quadratic Program (QP) at each sampling step. As the tube is loosest when noise is high and intervention is cheapest in terms of control energy, most constraint enforcement occurs when it least disrupts the model’s learned structure. We prove that this mechanism guarantees safe sampling while minimizing the distributional shift from the original model at each sampling step, as quantified by the KL divergence. Our framework applies to any pre-trained flow-based sampling scheme requiring no retraining or architectural modifications. We validate the approach across constrained image generation, physically-consistent trajectory sampling, and safe robotic manipulation policies, achieving 100% constraint satisfaction while preserving semantic fidelity.
@@ 4 Guidance of flow-based models via constricting CBFs
Having established the sampling dynamics (1) and the CBF framework for safe control, we address the central question: how can we enforce hard constraints on the output of a pre-trained generative model without modifying the model itself? The key observation is that by introducing a control input to the sampling SDE (1), we arrive at a control-affine structure for the control systems in Section 3.2. This additive control input $u$ can be designed to steer sampling trajectories toward the safe set while preserving the learned distribution. This is a control synthesis problem with the following guided sampling process:
@@ 4.1 CBFs for flow-based generative sampling
A fundamental challenge in generative sampling is that the initial prior noise $x(T)$ is drawn from a distribution such as the standard Gaussian. This means that $x(T)$ almost certainly resides outside the target safe set $\mathcal{C}$. Traditional CBF formulations for static sets would require an immediate, high-energy intervention to push $x(t)$ close to $\mathcal{C}$, which would violate the model’s generative intent and produce samples that no longer represent the learned distribution. Our framework is motivated by the need for minimal interventions and a smooth approach towards $\mathcal{C}$. We achieve this by defining a constricting barrier $\tilde{h}(x,t)=h(x)+\epsilon(x(T),t)$. This is a time-varying barrier that constricts with the sampling process. We characterize safety through a constricting superlevel tube $\tilde{\mathcal{C}}(t)$ with respect to the constricting barrier $\tilde{h}$ as,
@@
We design $\tilde{\mathcal{C}}(t)$ such that initially, the set $\tilde{\mathcal{C}}(T)$ is sufficiently relaxed to contain the noise sample $x(T)$, and at the end of sampling, it recovers the target set $\mathcal{C}$, i.e., $\tilde{\mathcal{C}}(0)=\mathcal{C}$. Since the sampling process evolves in reverse time from $t=T$ to 0, the safety tube $\tilde{\mathcal{C}}(t)$ must remain occupied by $x(t)$ as $t$ decreases, a property we term reverse invariance. As derived in Section 3.2, the reverse-time CBF condition requires bounding $\nabla\tilde{h}\cdot f$ from above. We design the constricting barrier function through a time-varying control barrier function.
@@
Given a CBF $h:\mathbb{R}^{n}\to\mathbb{R}$ and a terminal condition $x(T)$, a constricting barrier function is $\tilde{h}(x,t)=h(x)+\epsilon(x(T),t)$, where $\epsilon$ is a $C^{1}$ function satisfying:
@@
Unlike works that consider designing feedback controllers that enforce safety of stochastic dynamical systems (Clark, 2021), our approach assumes that the feedback controller has knowledge of the state as well as the noise. While this assumption is more restrictive, it allows us to obtain stronger guarantees. In particular, with our design approach, we can enforce invariance of the constricting safety tube exactly, whereas by So et al. (2023), invariance does not exactly or almost surely for the CBFs considered in Clark (2021).
@@
where the optimization is performed for the current state-noise pair $(x,\xi)$ at time $t$, yielding $u^{*}(x,\xi,t)$.
@@
$D_{\mathrm{KL}}(p_{\text{safe}}(0)\parallel p(0))=\frac{1}{2}\mathbb{E}_{p_{\text{safe}}(0)}\left[\int_{0}^{T}\frac{\|u(x,t)\|^{2}}{g(t)^{2}}\mathrm{d}t\right].$ (9)
@@ 4.2 Discrete-time implementation
We note that the discrete-time constraint (11) is derived via a first-order Taylor series expansion of $\tilde{h}$ around $(x_{k},t_{k})$, and is therefore only a sufficient condition for $\tilde{h}(x_{k-1},t_{k-1})\geq 0$ only up to linearization error. For barrier functions with high curvature or large step sizes $\Delta t$, the true nonlinear barrier value could deviate from the linear prediction. In practice, the margin provided by the class-$\mathcal{K}$ term $\gamma(\tilde{h}_{k})\Delta t$ and the conservative initialization $\epsilon_{0}$ absorb this error. Across all experiments in Section 5, we observed no constraint violations at the final sample, empirically validating the linearization for the step sizes used. The discrete-time constraint (11) provides an approximate safety guarantee: the first-order Taylor expansion introduces an $O(\Delta t^{2})$ residual that is not formally bounded in our analysis. An additive margin $c>0$ can be introduced in the initialization $\epsilon_{0}$ and the class-$\mathcal{K}$ damping $\gamma$ to provide empirical robustness to this discretization error, as well as provide a margin so that the optimization does not start from the zero-level set for $\tilde{h}(x,t)$. A rigorous discrete-time CBF analysis that establishes explicit step-size-dependent bounds on constraint violation is an important direction for future work. Nonetheless, across all experiments with step sizes ranging from $\Delta t$ = 0.01 (Section 5.1) to $\Delta t=0.005$ (Section 5.3), we observed zero constraint violations at the final sample.
@@
1. Initialize: Sample $x_{K}\sim p(T)$ (e.g., $x_{K}\sim\mathcal{N}(0,I)$), set $t_{K}=T$, $\Delta t=T/K$
2. for $k=K,K-1,\ldots,1$ do
 3. Sample noise: $\xi_{k}\sim\mathcal{N}(0,I)$
 4. Compute noise-induced drift: $d_{\text{noise}}=g(t_{k})\sqrt{\Delta t}\,\xi_{k}$
 5. Compute constricting barrier: $\tilde{h}_{k}=h(x_{k})+\epsilon(x_{K},t_{k})$
 6. Solve safety-constrained QP: $\min_{u_{k}}\quad$ $\frac{1}{2}\|u_{k}\|^{2}$ s.t. $\tilde{h}_{k}+\nabla_{x}\tilde{h}_{k}\cdot\left(-[f_{\theta}(x_{k},t_{k})+u_{k}]\Delta t+d_{\text{noise}}\right)-\frac{\partial\epsilon}{\partial t}(x_{K},t_{k})\Delta t\geq(1-\gamma(\tilde{h}_{k})\Delta t)\,\tilde{h}_{k}$
  0.5pt
 7. Apply control and update state: $x_{k-1}=x_{k}-f_{\theta}(x_{k},t_{k})\Delta t-u_{k}^{*}\Delta t+d_{\text{noise}}$
 8. Update time: $t_{k-1}=t_{k}-\Delta t$
9. return $x_{0}$
@@ 5 Experiments
All experiments use the linear constriction $\epsilon(x(T),t)=\epsilon_{0}\cdot\frac{t}{T}$ where $\epsilon_{0}=\max(0,-h(x(T)))+c$, where $c>0$ is a small positive margin. This margin prevents the barrier from operating exactly on its zero level set at initialization, where numerical precision could cause spurious violations. In all experiments, we set $c=0.1$. We choose the class-$\mathcal{K}$ function $\gamma(h)=\alpha h$ with $\alpha=0.5$. The QPs in Algorithm 1 are solved using CVXPY with the OSQP solver. Our implementation is available at [anonymized for review]. Our experiments were conducted on an Intel i9-9900 machine with 128GB RAM and an Nvidia Quadro RTX 4000 GPU.
@@ 5.1 Physics-consistent trajectory generation for the Lorenz system
Figure 2(b) reveals the evolution of the safety tube. At $t=1$, the noise sample $x(T)\sim\mathcal{N}(0,I)$ has large cumulative physics error, requiring relaxation $\epsilon_{0}\approx 16$—over four orders of magnitude larger than our target tolerance $e=0.001$. The blue line shows the linear constriction $\epsilon(x(T),t)=\epsilon_{0}\cdot(t/T)$ to zero at $t=0$. The constriction barrier $\tilde{h}(x,t)$ (red) remains non-negative throughout, verifying reverse invariance (Theorem 4.1). Fluctuations during $t\approx 0.8$-$1.0$ reflect the high-noise regime where $g(t)$ is large. At $t=0$, $\tilde{h}(x,0)\approx 0$ with $\epsilon(x(T),0)=0$ implies $h(x(0))\approx 0$, confirming that our minimum-norm control maximally exploits the constraint tolerance. Figure 2(c) shows control effort $\|u\|^{2}$ concentrated at sampling onset ($t\approx 0.8$-$1.0$), peaking around 130-140, then rapidly decaying to near-zero by $t\approx 0.6$. This front-loaded intervention exploits that control is cheaper when $g(t)$ is large. The minimal effort for $t<0.6$ indicates that the pre-trained model has implicitly learned the Lorenz system structure from training data, requiring only minor corrections to enforce hard guarantees.
@@ 5.2.2 Regional color intensity constraints
Figure 4: Regional color intensity constraints. Images are enforced with color values in the lower third. (a) Moderate intensity: The target pixel color is black, and the spatial mask $v(\mathbf{p})$ goes from $0.5\to 0$, allowing the diffusion model to produce clean furniture. (b) Projection: Projection-based constraint enforcement (Zampini et al., 2025), constraints are the same as in (a), with the target pixel color black in the bottom one-third of the image. Projection causes intense constraint enforcement, which leads to loss of semantic information. (c) Low intensity: Target color is brown, and the spatial mask $v(\mathbf{p})$ goes from $0.2$ at the bottom to 0 at the lower-third boundary, producing a semantically meaningful image of a room with a brown carpet.
@@ 5.3 Smooth robot policy generation
where $e_{\text{smooth}}=1.5$ is the tolerance that bounds the maximum allowed jerk magnitude, and $\Delta_{s}$ is the time discretization of the environment. The safe set $\mathcal{C}=\{x\mid h(x)\geq 0\}$ encodes the set of all action sequences whose average control variation magnitude remains below the prescribed threshold $e_{\text{smooth}}$. We apply our Algorithm 1 with the same linear constriction scheme as in the previous experiments.
@@
Table 1: Comparison on Push-T with smoothness constraints over 100 episodes
@@ 6 Limitations and future work
Several extensions follow naturally from our framework. First, extending the approach to latent diffusion models would enable scalability to higher-resolution generation. We investigated extending our framework to latent diffusion models (Rombach et al., 2022) where the diffusion process operates in a compressed latent space $z=E(x)$ and the constraint is defined in image space. While the QP structure is preserved, the latent gradient is computed via a single backward pass through the decoder. Our preliminary experiments with Stable Diffusion v1.5 revealed that exact constraint satisfaction does not transfer from latent to image space (see Appendix B.2). The VAE decoder is not a diffeomorphism: it is locally non-invertible and introduces reconstruction artifacts that distort the constraint boundary. As a result, the CBF shield successfully biases the latent trajectory toward the constrained region but cannot guarantee pixel-level fidelity in the decoded image. Achieving exact constraints in latent diffusion models likely requires either (i) a constraint-aware fine-tuning of the decoder to improve local invertibility near the constraint boundary, or (ii) a hybrid approach that applies coarse guidance in latent space and a final correction step in pixel space after decoding. Second, enforcing state-space safety constraints in robotic policies, such as collision avoidance, requires a dynamics model to map actions to states. Combining our framework with learned dynamics models or neural ODEs is a promising direction. Third, our current implementation solves one QP per sampling step using a general-purpose solver. Using an MPC framework across many steps could significantly yield tighter bounds on the KL divergence over the entire sampling horizon. In particular, a receding-horizon formulation that anticipates the constriction schedule over multiple future steps could yield control policies that are globally optimal, rather than greedy at each step, potentially tightening the bound over the learned sampling process.